OpenStudy (anonymous):

Last year 4 local bands played at Beach High School’s spring break concert. Each band was paid $100, and the remaining $2725 in ticket sales went to the school’s recreation fund. If 6 bands will have to be paid $100 each this year and tickets are still $5 each, Q: How many tickets will have to be sold this year in order to raise the same amount of money as last year for the school’s recreation fund? ~please i want to know how to find the answer show show your work~

OpenStudy (anonymous):

Enough cash would need to be raised to cover the 600 that will be paid to bands and the 2725 that we need for the recreation fund. so we need to raise 600+2725=3325... since tickets are 5 each, number of tickets that need to be sold = (3325/5)=665

OpenStudy (anonymous):

wow thx KBOY i was distracted by something else--word problems are tough for me sometimes